OpenStudy (anonymous):

Write a radical in which the radicand consists of a coefficient and a variable. Use a coefficient that can be simplified best using the greatest perfect square method. Then, simplify the radical completely. You can use "sqrt()" to show a square root. For example, 7 times the square root of a can be written as 7sqrt(a).

OpenStudy (anonymous):

hold on ill help u :3

OpenStudy (anonymous):

hope that this will help you understanding - so radical indice 2 sign squareroot what is equal in exponentialy form by 1/2 - radical indice 3 is cuberoot what in exponent form is 1/3 - so for your help for example 4^(1/2 = (2^2)^(1/2) =2^(2*1/2) = =2^(2/2) =2^1 =2 - but this is squareroot of 4 =2 - ok ?

OpenStudy (anonymous):

4? is that right? 4 x 4 = 16, and 16 x 4 = 64

OpenStudy (anonymous):

I believe so

OpenStudy (anonymous):

i had a lil trouble with this so im not entirely sure

OpenStudy (anonymous):

there you go :)

OpenStudy (anonymous):

i searched it so i know its right and i asked my teacher lol

OpenStudy (anonymous):

ty so much